OpenStudy (brn_betty):

Differentiate. f(x) = 5 x sin x

OpenStudy (brn_betty):

\[5\sqrt{x}\sin x\]

jimthompson5910 (jim_thompson5910):

use the product rule let f(x) = sqrt(x) and g(x) = sin(x)

jimthompson5910 (jim_thompson5910):

the 5 is a coefficient which can be temporarily pulled out, and ignored (but don't forget to place it back in when you're done deriving)

OpenStudy (brn_betty):

ok thats what i was doing wrong. Forgot to put the 5 back in. Thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (brn_betty):

Find the derivative of the function. y=3^(8-x^2)

jimthompson5910 (jim_thompson5910):

you will use the exponential rule http://www.mathsisfun.com/calculus/derivatives-rules.html and the chain rule

jimthompson5910 (jim_thompson5910):

when you get to 8-x^2, you'll use the power rule

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (brn_betty):

-2x(ln3)3^(8-x^2)

OpenStudy (brn_betty):

does thae same rule apply to base g? or is that a constant? as in \[g^{\sin pix}\]

OpenStudy (brn_betty):

oh never mind it actually says 9 not g.

jimthompson5910 (jim_thompson5910):

-2x(ln3)3^(8-x^2) is correct

jimthompson5910 (jim_thompson5910):

and if they say "derive with respect to x", then the other letters are usually constants (unless they state something like "g is a function of x")

OpenStudy (brn_betty):

A stone is dropped into a lake, creating a circular ripple that travels outward at a speed of 50 cm/s. Find the rate at which the area within the circle is increasing after 3 seconds. I get 47,123 instead of 15000pi. What am i missing

jimthompson5910 (jim_thompson5910):

"A stone is dropped into a lake, creating a circular ripple that travels outward at a speed of 50 cm/s". So this means that the radius is growing at a speed of 50 cm/s. So let's first find the radius function (function of time) Length of Radius = (Speed)*(Time) r = s*t \[\Large r = 50t\] Then derive this radius function with respect to t. This is effectively the same as its speed. \[\Large \frac{dr}{dt} = 50\] ------------------------------------------------------- Plug t = 3 into the equation r = 50t to get the radius at time t = 3 \[\Large r = 50t\] \[\Large r = 50(3)\] \[\Large r = 150\] At time t = 3, the radius is 150 cm. ------------------------------------------------------- Now derive the area function with respect to t and plug in the values found above \[\Large A = \pi*r^2\] \[\Large \frac{dA}{dt} = 2\pi*r*\frac{dr}{dt}\] \[\Large \frac{dA}{dt} = 2\pi*150*50\] \[\Large \frac{dA}{dt} = (2*150*50)*\pi\] \[\Large \frac{dA}{dt} = 15,000\pi\]